A market's demand and supply are given as follows: Supply is equal to 3p, and demand is equal to 120-5p. An excise tax is imposed on this market at a rate of $8 per traded unit. At equilibrium, what is value of the tax revenue?
1
Expert's answer
2017-10-09T15:17:06-0400
Equilibrium price is:
3p=120-5p
8p=120
P=$15
Equilibrium quantity is:
Q=3*15 = 45 or 120-5*15 = 45
Value of the tax revenue is:
T=45*$8=$360